OpenStudy (anonymous):

What is the quotient: (3x3 – 16x2 + 23x – 6) ÷ (3x – 1) ?

OpenStudy (anonymous):

long division ( a pain) or cheating

OpenStudy (anonymous):

Idkhow to do this.

OpenStudy (anonymous):

turns out the numerator factors as \[(3x-1)(x-3)(x-2)\]

OpenStudy (anonymous):

Well, you want to look at it as if it's just any other long division problem. (3x - 1) * (-x^2) => (-3x^3 + x^2) (–3x3 + 16x2 + 13x – 6) - (-3x^3 + x^2) = (15x^2 + 13x - 6) (3x - 1) * (5x) => (15x^2 - 5x) (15x^2 + 13x - 6) - (15x^2 - 5x) = (18x - 6) (3x - 1) * 6 = 18x - 6 (18x - 6) - (18x - 6) = 0 So, the final result will be => -x^2 + 5x + 6

OpenStudy (anonymous):

here is the answer, but long division is a pain http://www.wolframalpha.com/input/?i=%283x3+%E2%80%93+16x2+%2B+23x+%E2%80%93+6%29+%2F+%283x+%E2%80%93+1%29+